# Photodynamic melting of phase-reversed charge stripes and enhanced condensation

### v1

@@ 2 Relevant Matrix elements
The perturbative analysis above emphasizes the role of the current and kinetic energy operators in the fate of dynamically accessing target states. In the main text, we have focused on the $\hat{J}_{x}$ matrix elements in the eigenbasis of the unperturbed Hamiltonian. Now, we extend this analysis by showing the corresponding $y$-direction in Fig. S1(b). Unlike the $x$-direction case, the matrix $\langle m|\hat{J}_{y}|n\rangle$ is much sparser and does not exhibit significant matrix elements of the form $\langle 2|\hat{J}_{y}|0\rangle$, thus precluding first-order excitation of the target state $|2\rangle$, nor simultaneously large $\langle 2|\hat{J}_{y}|m^{\prime}\rangle$ and $\langle m^{\prime}|\hat{J}_{y}|0\rangle$, preventing relevant second-order current-current excitation terms. This justifies the choice of $x$-polarized pulses we implement in the main text.
@@
Figure S1: The absolute value of the current matrix elements in the eigenstate basis of the Hamiltonian, $|\langle m|\hat{J}_{\mu}|n\rangle|$, focusing on the 30 states with lowest energy, for $\mu=x$ in (a) and $\mu=y$ in (b). The dashed lines highlight the target state $|2\rangle$ used in the main text.
@@
Additionally, a strong indication that a second-order process is responsible for the excitation of eigenstate $|2\rangle$ is given by the analysis of the matrix elements of $\hat{J}_{x}^{2}$ in the eigenbasis of $\hat{\cal H}$. As Eqs. (S12) and (S24) indicate, relevant terms as $\propto\sum_{m^{\prime}}\langle m|\hat{J}_{x}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}|0\rangle=\langle m|\hat{J}_{x}^{2}|0\rangle$ need to be large for this type of contribution to be important. Indeed, as Fig. S2(b) shows, the most significant matrix elements of $\hat{J}_{x}^{2}$ in the low-lying spectrum are in between states $|2\rangle$ and $|0\rangle$.
@@
Finally, contributions for a first-order transition between $|0\rangle$ and $|2\rangle$ in the time-dependent perturbation can be accomplished by using the kinetic energy operator, see Eq. (S22), with matrix elements proportional to $\langle m|\hat{K}_{x}|0\rangle$. As Fig. S2(a) indicates, $|\langle 2|\hat{K}_{x}|0\rangle|$ is the largest in the low-lying spectrum. Nonetheless, because the coefficient in the first-order transition amplitude is $A^{2}(t)$ for the kinetic energy operator, unlike $A(t)$ for the current one, this dictates that such kinetic-driven excitation is not significantly favored for small pulse amplitudes.
@@
Figure S2: The absolute value of the matrix elements of the $x$-direction kinetic energy (a) and $x$-direction squared-current (b) operators in the eigenstate basis of the Hamiltonian, focusing on the 12 states with lowest energy. Here, the diagonal elements are set to zero for visualization (they are typically much larger). The dashed lines highlight the target state $|2\rangle$ used in the main text.
3 Density dynamics
@@
The main text focuses on the melting of the charge order reversal with the pulses, releasing a suppressed zero-momentum occupancy. Meanwhile, we can also investigate the fate of the imposed stripes by the local chemical potential $V_{0}$ under photoirradiation. Figure S3 shows the dynamics of the densities for the four inequivalent sites (under periodic boundary conditions) through photoirradiation with the same conditions as Figs. 1 and 3 of the main text, i.e., under maximization of the late-time overlap $|\langle 2|\psi(t_{\rm max})\rangle|$.
@@
Figure S3: Dynamics for the density resolved in the inequivalent sites under periodic boundary conditions – inset depicts the lattice with inequivalent sites annotated from 0 to 3. Photoirradiation conditions are the same of Fig. 1 in the main text: $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$, for Hamiltonian parameters $V/t_{h}=4$ and $V_{0}/t_{h}=5$.
@@
Figure S4: Dependence of the long-time average of the zero momentum occupancy in the pulse parameters $A_{0}$, pulse amplitude, and $t_{d}$, pulse width. Here the pulse frequency is the one used in Figs. 1 and 3 of the main text, $\Omega=1.713\ t_{h}$. The white-edge marker describes the pair of parameter values ($A_{0},t_{d}$) used in the main text.
4 Maximizing zero-momentum occupancy
@@
In the main text, we establish what pump parameters, namely $A_{0}$, $t_{d}$, and $\Omega$, maximize the overlap with the target state, the equilibrium eigenstate $|2\rangle$ of $\hat{\cal H}$, as emphasized in Fig. 4(a). We now show that this set of parameters also maximizes the late-time averaged zero momentum occupancy $\langle\hat{n}_{{\bf k}=0}\rangle$ sufficiently after the photoirradiation. Figure S4 shows the time average of the zero momentum occupancy at long times ($t\in[4.5t_{d}-4t_{h}^{-1},4.5t_{d}]$) as a function of the pump amplitude and width for the frequency $\Omega=1.713\ t_{h}$ — it is clear that condensation is enhanced in the same range of parameters in which the overlap $|\langle\psi(t\gg t_{d})|2\rangle|^{2}$ is maximized.
@@
Figure S5: Dependence on $A_{0}$ and $t_{d}$ of the long-time weights of equilibrium states in $|\psi(t)\rangle$, for $|m=0\rangle$ (a), $|2\rangle$ (b), and $|4\rangle$ (c) — the marker assigns the parameters $(A_{0},t_{d})=(0.34,10.8)$ resulting in the maximum overlap with the target state $|2\rangle$. (d) Pulse representation for these parameter settings. Dynamics of the connected density correlations across a stripe (e), and the zero momentum occupancy (f). (g) Selected overlaps, by higher values in the dynamics, of the time-evolving state during the photoirradiation process. Data in the plots employ the second-order resonant photoirradiation $\Omega=1.491\ t_{h}=\omega_{42}$.
5 Other intermediate states
@@
While analyzing the matrix elements of the current operator $\hat{J}_{x}$ [Fig. S1], we notice that other states can, in principle, serve as intermediate states for a dynamical transition $|0\rangle\to|2\rangle$. For instance, considering $|m^{\prime}=4\rangle$, we notice that $|\langle 2|\hat{J}_{x}|4\rangle|=0.24$ and $|\langle 4|\hat{J}_{x}|0\rangle|=1.73$ — the largest among the lowest 30 eigenpairs of $\hat{\cal H}$. Thus, using a similar prescription of the main text for the case $|m^{\prime}=6\rangle$, resonant and energy conservation conditions lead, respectively, to $E_{4}-E_{0}\approx n_{1}\Omega$ and $E_{4}-E_{0}=(n_{1}+n_{2})\Omega$. Subtracting these, we obtain $\omega_{42}\simeq-n_{2}\Omega$. As in the main text, if we consider the $|4\rangle\to|2\rangle$ dynamical process as a one-photon emission ($n_{2}=-1$), we can set the pump frequency $\Omega=\omega_{42}\simeq 1.491$. Now combining with the resonant condition, yields $\omega_{40}/\Omega\approx n_{1}=4.475/1.491\simeq 3$, indicating a similar three quantum absorption scenario.
@@
Besides that, additional tuning of the pulse amplitudes and time-width is needed in order to target the desired state, $|2\rangle$, properly. This is accomplished by analyzing the long-time weight of selected equilibrium eigenstates $|m\rangle$ in $|\psi(t)\rangle$, as shown in Fig. S5(a)–(c). Maximimization of $|\langle 2|\psi(t_{\rm max})\rangle|^{2}$ is achieved, with $\Omega=\omega_{42}=1.491\ t_{h}$ at $A_{0}=0.34$ and $t_{d}=10.8\ t_{h}^{-1}$. Focusing on these parameters [see Fig. S5(d) for the pulse representation], we show the dynamics of both the connected density correlations [Fig. S5(e)] as well as the zero momentum occupancy [Fig. S5(f)]. In this case, the enhancement of the average $n_{{\bf k}=0}$ of about $26\%$ after the pulse is masked by the substantial oscillations that well encompass values below that of the equilibrium regime. Moreover, despite the reduction of the $\pi$-phase density correlations across a stripe, their melting is not achieved in these settings. This suggests that the dynamical addressing of the target state was not optimized.
@@
Indeed, as the dynamics of the overlaps in Fig. S5(g) displays, the weight of $|2\rangle$ in $|\psi(t)\rangle$ is yet small in comparison with the one with the initial state, the ground-state of $|0\rangle$. As a consequence, the targeting is only partial, in contrast to that obtained in the main text.

### v2

@@ Appendix B Relevant Matrix elements
The perturbative analysis above emphasizes the role of the current and kinetic energy operators in the fate of dynamically accessing target states. In the main text, we have focused on the $\hat{J}_{x}$ matrix elements in the eigenbasis of the unperturbed Hamiltonian. Now, we extend this analysis by showing the corresponding $y$-direction in Fig. 6(b). Unlike the $x$-direction case, the matrix $\langle m|\hat{J}_{y}|n\rangle$ is much sparser and does not exhibit significant matrix elements of the form $\langle 2|\hat{J}_{y}|0\rangle$, thus precluding first-order excitation of the target state $|2\rangle$, nor simultaneously large $\langle 2|\hat{J}_{y}|m^{\prime}\rangle$ and $\langle m^{\prime}|\hat{J}_{y}|0\rangle$, preventing relevant second-order current-current excitation terms. This justifies the choice of $x$-polarized pulses we implement in the main text.
@@
Figure 6: The absolute value of the current matrix elements in the eigenstate basis of the Hamiltonian, $|\langle m|\hat{J}_{\mu}|n\rangle|$, focusing on the 30 states with lowest energy, for $\mu=x$ in (a) and $\mu=y$ in (b). The dashed lines highlight the target state $|2\rangle$ used in the main text.
@@
Additionally, a strong indication that a second-order process is responsible for the excitation of eigenstate $|2\rangle$ is given by the analysis of the matrix elements of $\hat{J}_{x}^{2}$ in the eigenbasis of $\hat{\cal H}$. As Eqs. (15) and (27) indicate, relevant terms as $\propto\sum_{m^{\prime}}\langle m|\hat{J}_{x}|m^{\prime}\rangle\langle m^{\prime}|\hat{J}_{x}|0\rangle=\langle m|\hat{J}_{x}^{2}|0\rangle$ need to be large for this type of contribution to be important. Indeed, as Fig. 7(b) shows, the most significant matrix elements of $\hat{J}_{x}^{2}$ in the low-lying spectrum are in between states $|2\rangle$ and $|0\rangle$.
@@
Finally, contributions to a first-order transition between $|0\rangle$ and $|2\rangle$ under the time-dependent perturbation can be obtained using the kinetic energy operator, see Eq. (25), with matrix elements proportional to $\langle m|\hat{K}_{x}|0\rangle$. As Fig. 7(a) indicates, $|\langle 2|\hat{K}_{x}|0\rangle|$ is the largest in the low-lying spectrum. Nonetheless, because the coefficient in the first-order transition amplitude is $A^{2}(t)$ for the kinetic energy operator, unlike $A(t)$ for the current one, this dictates that such kinetic-driven excitation is not significantly favored for small pulse amplitudes.
@@
Figure 7: The absolute value of the matrix elements of the $x$-direction kinetic energy (a) and $x$-direction squared-current (b) operators in the eigenstate basis of the Hamiltonian, focusing on the 12 states with lowest energy. Here, the diagonal elements are set to zero for visualization (they are typically much larger). The dashed lines highlight the target state $|2\rangle$ used in the main text.
Appendix C Density dynamics
@@
The main text focuses on the melting of the charge-order reversal under the pulses, releasing the suppressed zero-momentum occupancy. Meanwhile, we can also investigate the fate of the imposed stripes under photoirradiation in terms of the local chemical potential $V_{0}$. Figure 8 shows the dynamics of the densities for the four inequivalent sites (under periodic boundary conditions) through photoirradiation with the same conditions as Figs. 1 and 3 of the main text, i.e., under maximization of the late-time overlap $|\langle 2|\psi(t_{\rm max})\rangle|$.
@@
Figure 8: Dynamics for the density resolved in the inequivalent sites under periodic boundary conditions – inset depicts the lattice with inequivalent sites annotated from 0 to 3. Photoirradiation conditions are the same of Fig. 1 in the main text: $\Omega=1.713\ t_{h}$, $A_{0}=0.62$ and $t_{d}=6.5t_{h}^{-1}$, for Hamiltonian parameters $V/t_{h}=4$ and $V_{0}/t_{h}=5$.
@@
Figure 9: Dependence of the long-time average of the zero momentum occupancy in the pulse parameters $A_{0}$, pulse amplitude, and $t_{d}$, pulse width. Here, the pulse frequency is the same as in Figs. 1 and 3 of the main text, $\Omega=1.713\ t_{h}$. The white-edge marker describes the pair of parameter values ($A_{0},t_{d}$) used in the main text.
Appendix D Maximizing zero-momentum occupancy
@@
In the main text, we establish what pump parameters, namely $A_{0}$, $t_{d}$, and $\Omega$, maximize the overlap with the target state, the equilibrium eigenstate $|2\rangle$ of $\hat{\cal H}$, as emphasized in Fig. 4(a). We now show that this set of parameters also maximizes the late-time averaged zero momentum occupancy $\langle\hat{n}_{{\bf k}=0}\rangle$ sufficiently after the photoirradiation. Figure 9 shows the time average of the zero momentum occupancy at long times ($t\in[4.5t_{d}-4t_{h}^{-1},4.5t_{d}]$) as a function of the pump amplitude and width for the frequency $\Omega=1.713\ t_{h}$ — it is clear that condensation is enhanced in the same range of parameters in which the overlap $|\langle\psi(t\gg t_{d})|2\rangle|^{2}$ is maximized.
@@
Figure 10: Dependence on $A_{0}$ and $t_{d}$ of the long-time weights of equilibrium states in $|\psi(t)\rangle$, for $|m=0\rangle$ (a), $|2\rangle$ (b), and $|4\rangle$ (c) — the marker assigns the parameters $(A_{0},t_{d})=(0.34,10.8)$ resulting in the maximum overlap with the target state $|2\rangle$. (d) Pulse representation for these parameter settings. Dynamics of the connected density correlations across a stripe (e), and the zero momentum occupancy (f). (g) Selected overlaps, by higher values in the dynamics, of the time-evolving state during the photoirradiation process. Data in the plots employ the second-order resonant photoirradiation $\Omega=1.491\ t_{h}=\omega_{42}$.
Appendix E Other intermediate states
@@
While analyzing the matrix elements of the current operator $\hat{J}_{x}$ [Fig. 6], we notice that other states can, in principle, serve as intermediate states for a dynamical transition $|0\rangle\to|2\rangle$. For instance, considering $|m^{\prime}=4\rangle$, we notice that $|\langle 2|\hat{J}_{x}|4\rangle|=0.24$ and $|\langle 4|\hat{J}_{x}|0\rangle|=1.73$ — the largest among the lowest 30 eigenpairs of $\hat{\cal H}$. Thus, using a similar prescription of the main text for the case $|m^{\prime}=6\rangle$, resonant and energy conservation conditions lead, respectively, to $E_{4}-E_{0}\approx n_{1}\Omega$ and $E_{4}-E_{0}=(n_{1}+n_{2})\Omega$. Subtracting these, we obtain $\omega_{42}\simeq-n_{2}\Omega$. As in the main text, if we consider the $|4\rangle\to|2\rangle$ dynamical process as a one-photon emission ($n_{2}=-1$), we can set the pump frequency $\Omega=\omega_{42}\simeq 1.491$. Now combining with the resonant condition, yields $\omega_{40}/\Omega\approx n_{1}=4.475/1.491\simeq 3$, indicating a similar three quantum absorption scenario.
@@
Besides that, additional tuning of the pulse amplitudes and time-width is needed in order to target the desired state, $|2\rangle$, properly. This is accomplished by analyzing the long-time weight of selected equilibrium eigenstates $|m\rangle$ in $|\psi(t)\rangle$, as shown in Fig. 10(a)–(c). Maximimization of $|\langle 2|\psi(t_{\rm max})\rangle|^{2}$ is achieved, with $\Omega=\omega_{42}=1.491\ t_{h}$ at $A_{0}=0.34$ and $t_{d}=10.8\ t_{h}^{-1}$. Focusing on these parameters [see Fig. 10(d) for the pulse representation], we show the dynamics of both the connected density correlations [Fig. 10(e)] as well as the zero momentum occupancy [Fig. 10(f)]. In this case, the enhancement of the average $n_{{\bf k}=0}$ of about $26\%$ after the pulse is masked by the substantial oscillations that well encompass values below that of the equilibrium regime. Moreover, despite the reduction of the $\pi$-phase density correlations across a stripe, their melting is not achieved in these settings. This suggests that the dynamical addressing of the target state was not optimized.
@@
Indeed, as the dynamics of the overlaps in Fig. 10(g) displays, the weight of $|2\rangle$ in $|\psi(t)\rangle$ is yet small in comparison with the one with the initial state, the ground-state of $|0\rangle$. As a consequence, the targeting is only partial, in contrast to that obtained in the main text.
A critical discussion we left aside in our analysis concerns the system size dependence of our results. As we establish, resonant conditions depend on the low-lying spectrum properties of the relevant Hamiltonians. The issue is that, with larger system sizes, while the energy levels (such as those for the ground state and the target state) are extensive quantities that significantly change with the system size considered, their difference (i.e., the gaps) required to find resonance conditions via the pump has only subextensive corrections. This means that the results could, in principle, be generalized to larger system sizes without the need for dramatic fine-tuning of the pump frequencies when approaching the thermodynamic limit, thereby attesting to the generality of our findings.
Appendix F Decay of Single-Particle Correlations
In the main text, we use quantities to probe condensation and phase coherence, the zero-momentum occupancy $\hat{n}_{{\bf k}=0}$ and the condensate fraction $f_{0}$, which inherently build on top of the single-particle correlations, forming the single-particle density matrix (SPDM) correlations,
$C(r,t)=\langle\hat{b}_{{\bf i}}^{\dagger}(t)\hat{b}_{\bf i+r}^{\phantom{\dagger}}(t)\rangle\ .$ (30)
[…]
The crucial and interesting point is that, at long distances, the correlations at late time ($t>0$) do not follow a simple power-law behavior. Instead, $C(r,t)/C(r=1,t)$ tends to saturate toward a finite value, signaling the emergence of a nonvanishing off-diagonal component of the single-particle density matrix. This plateau indicates that the coherence length $\xi(t)$ exceeds the system size, corresponding to the development of global phase coherence and long-range order. In contrast, before the pulse ($t<0$), the correlations decay rapidly, consistent with a short-range correlated stripe state. The observed saturation at long times thus represents the finite-size manifestation of photoinduced condensation, in agreement with the enhanced condensate fraction $f_{0}(t)$ and Drude weight reported in the main text.
Figure 11: Spatial decay of single-particle correlations $C(r,t)/C(r=1,t)$ at representative times indicated by the vertical arrows in the pump profile (upper panel). Warm (cool) colors correspond to pre- (post-) irradiation states. After the pulse, the correlations saturate at large distances, signaling the emergence of long-range coherence consistent with the enhanced condensate fraction $f_{0}(t)$ reported in the main text. The inset displays the same data on a log-log scale. Here, the pump parameters are the same as in Fig. 1 of the main text.
Appendix G Distinction between Drude and superfluid weights
In the main text, we introduce a protocol for the out-of-equilibrium determination of the Drude weight $D$, based on the insertion of a probe linear-in-time vector potential much after the photoirradiation, when the system is closely relaxing toward equilibrium conditions. We now further justify this approach and contrast it with one that enables the out-of-equilibrium determination of a metric of phase coherence, namely the superfluid weight $D_{s}$.
We start by reviewing the conditions under which one extracts the equilibrium $D$ and $D_{s}$. In equilibrium, the current response to a $x$-oriented vector potential $A_{x}$ can be expressed as [62]
$\langle\hat{J}_{x}\rangle=-\langle\hat{K}_{x}\rangle A_{x}+\langle\hat{\Pi}_{xx}({\bf q},\omega)\rangle A_{x}\,$, (31)
where $\langle\hat{K}_{x}\rangle$ and $\langle\hat{\Pi}_{xx}\rangle$ are the diamagnetic and paramagnetic contributions, respectively, the latter mapping the current-current correlator. Two distinct stiffnesses arise depending on the order of limits taken for $\omega$ and $\bf q$}:
$D^{(x)}$ $=\langle-\hat{K}_{x}\rangle-\lim_{\omega\to 0}\langle\hat{\Pi}_{xx}({\bf q}{=}0,\omega)\rangle,$
$D_{s}^{(x)}$ $=\langle-\hat{K}_{x}\rangle-\lim_{q_{x}=0,q_{y}\to 0}\langle\hat{\Pi}_{xx}({\bf q},\omega{=}0)\rangle.$ (32)
The first quantity, $D^{(x)}$, is the equilibrium Drude weight or charge stiffness, governing the zero-frequency weight of the optical conductivity and quantifying ballistic charge transport in the long-wavelength (${\bf q}=0$) limit. The second, $D_{s}^{(x)}$, is the equilibrium superfluid weight, obtained from the static ($\omega=0$) transverse limit ($q_{x}=0,q_{y}\!\to\!0$) and associated with phase rigidity and the Meissner response.
An equivalent, and often more transparent, way to express these stiffnesses is in terms of the curvature of the total energy with respect to a uniform twist $\Phi_{x}=A_{x}L_{x}$ applied along the periodic direction [61, 62]. For the ground state $|m=0\rangle$ with energy $E_{0}(\Phi_{x})$, one can write
$\frac{D^{(x)}}{\pi}=\frac{L_{x}^{2}}{L_{x}L_{y}}\left.\frac{\partial^{2}E_{0}(\Phi_{x})}{\partial\Phi_{x}^{2}}\right|_{\Phi_{x}=0}\,.$ (33)
In finite systems, $E_{0}(\Phi_{x})$ generally exhibits cusps whenever the ground-state wave function crosses another state with a different current quantum number. As discussed in Ref. [62], in one dimension, such crossings remain well-separated even in the thermodynamic limit, allowing one to adiabatically follow the ground state and unambiguously evaluate the curvature of $E_{0}(\Phi_{x})$ at $\Phi_{x}=0$, which yields the Drude weight. In higher dimensions, however, the crossings proliferate and become dense, so that the ground-state energy as a function of twist is instead represented by the smooth envelope of the many-body branches $E_{n}(\Phi_{x})$. In this case, computing $\partial^{2}E/\partial\Phi_{x}^{2}$ at fixed system size and extrapolating gives the adiabatic curvature associated with the Drude weight $D$, whereas taking the thermodynamic limit first and differentiating the envelope gives the superfluid weight $D_{s}$. The two quantities coincide in one-dimensional systems, but differ in higher dimensions or whenever level crossings destroy adiabatic continuity.
These definitions, expressed in terms of energy derivatives, are fully equivalent to those from the current response, provided the same limit ordering is respected. In equilibrium, the distinction between the Drude and superfluid weights stems from the different order in which temporal and spatial limits are taken: taking $\omega\to 0$ before ${\bf q}\to 0$ (the long wavelength limit) yields the Drude weight $D$, while reversing the order (the static transverse case) gives the superfluid weight $D_{s}$. This distinction can be dynamically mirrored by preparing nonequilibrium states that realize, respectively, a uniform time-dependent or a static spatially twisted gauge field.
To make this correspondence explicit, recall that in a gauge where the scalar potential is zero, the electric field follows from the time derivative of the vector potential, $E_{x}(t)=-\partial_{t}A_{x}(t)$. A slowly varying, linear-in-time ramp of $A_{x}(t)$ therefore produces a uniform, nearly static electric field and probes the small-frequency, spatially uniform response, i.e., the ${\bf q}=0$, $\omega\!\to\!0$ limit relevant to charge transport. In contrast, a static vector potential $\delta A_{x}$ has $\partial_{t}A_{x}=0$ and thus no electric field. Its effect enters solely through the spatial boundary conditions, or equivalently, through a uniform twist $\Phi_{x}=A_{x}L_{x}$ in the many-body wavefunction, which probes the system’s response to a static phase gradient. This corresponds to the static ($\omega=0$), long-wavelength (${\bf q}\!\to\!0$) limit that underlies the equilibrium definition of the superfluid weight.
Figure 12: (a) Time evolution of the current $\langle J_{x}(t)\rangle$ after static vector-potential quenches $A_{x}=\delta A_{x}$ applied at $t_{1}=5t_{d}$. The original photoirradiation pulse, centered at $t_{0}=0$, is shown at the top. (b) Late-time average $\overline{\langle J_{x}\rangle}$ [time-window given by the shaded region in (a)] versus $\delta A_{x}$, whose linear slope determines $D_{s}^{(\mathrm{eff},x)}=-\partial J_{x}/\partial A_{x}|_{A_{x}\to 0}$. Dashed line displays a linear fit to the data, fitting parameters in the legend, whereas error bars are the standard error of the mean of $\langle J_{x}(t)\rangle$ within the averaging window. Pump parameters are the same as Fig. 1 of the main text, and of Fig. 11.
In this way, the ramp and static protocols directly implement, in real time, the two complementary limits that distinguish $D$ and $D_{s}$ in equilibrium. The former probes the inertial response to an applied electric field (longitudinal conductivity), while the latter probes the phase rigidity of the current in a time-independent twist (transverse superfluid response). The resulting persistent current encodes the slope of the smooth envelope of the energy with respect to the twist, [62]
$D_{s}^{({\rm eff},x)}=-\frac{\partial J_{x}}{\partial A_{x}}\Big|_{A_{x}\to 0}\,$, (34)
providing a nonequilibrium analogue of the superfluid weight.
Thus, to quantify the out-of-equilibrium superfluid response, we promote vector-potential quenches $A_{x}=\delta A_{x}$ at times $t_{1}=5t_{d}$, where the pulse has significantly faded. Computing the current response $\langle J_{x}(t)\rangle$ for various $\delta A_{x}$ [Fig. 12(a)] allows one to extract $D_{s}^{(\mathrm{eff},x)}$ from the slope of the steady-state current $\overline{\langle J_{x}\rangle}$ versus $\delta A_{x}$ [Fig. 12(b)]. The relation $J_{x}=-D_{s}^{(x)}A_{x}$ implies that a positive superfluid stiffness corresponds to a restoring current flowing opposite to the imposed vector potential. The simultaneous finite values of $D^{(x)}$ (see main text) and $D_{s}^{(x)}$ thus demonstrate that photoirradiation induces not only ballistic charge transport but also a phase-coherent superfluid response.
Appendix H Symmetries and targeting of states
In the above discussion, we explore how time-dependent perturbation theory provides a framework for understanding which states can be targeted via matrix elements of relevant operators, including the current $\hat{J}_{x}$. In particular, because this operator is odd under reflection symmetry, a selection rule emerges, wherein only eigenstates with opposite parities may possess a finite matrix element, i.e., $\langle{\rm odd}|\hat{J}_{x}|{\rm even}\rangle\neq 0$, whereas $\langle{\rm even}|\hat{J}_{x}|{\rm even}\rangle$ and $\langle{\rm odd}|\hat{J}_{x}|{\rm odd}\rangle$ are identically zero.
This can be seen by considering the reflection operator $\hat{\cal R}_{3}$, a reflection about the stripe located at $x=3$ [see Fig. 1(a)], which, combined with the periodic boundary conditions, leads to the mapping of the site coordinates $\hat{\cal R}_{3}:(x,y)\mapsto(-x\mod L_{x},y)$. If labeling the sites by their $x$-coordinate in our $L_{x}=8$ ladder, this means this map preserves the stripe sites, $3\mapsto 3$ and $7\mapsto 7$, while symmetric sites about a stripe are mapped to one another: $2\mapsto 4$, $1\mapsto 5$, and $0\mapsto 6$. Since $\hat{\cal R}_{3},\hat{\cal H}]=0$, the eigenstates $|m\rangle$ of $\hat{\cal H}$ are also eigenstates of $\hat{\cal R}_{3}$ and possess a definite parity, $\hat{\cal R}_{3}|m\rangle=\pm|m\rangle$.
Turning back to the symmetry of the current operator, one sees that it is antisymmetric under the reflection $\hat{\cal R}_{3}$, i.e., $\hat{\cal R}_{3}\hat{J}_{x}\hat{\cal R}_{3}^{-1}=-\hat{J}_{x}$. As a consequence, the matrix elements obey
$\langle m|\hat{J}_{x}|n\rangle=-(\pm_{m})(\pm_{n})\langle m|\hat{J}_{x}|n\rangle,$
which can only be satisfied if the states $|m\rangle$ and $|n\rangle$ possess opposite parities. Inspecting Fig. 4(c), one therefore finds that the even-parity ground state $|0\rangle$ has finite current matrix elements only with odd-parity excited states, namely $|3\rangle$, $|4\rangle$, $|6\rangle$, and $|10\rangle$. By contrast, the remaining low-lying states $|1\rangle$, $|2\rangle$, $|5\rangle$, $|7\rangle$, $|8\rangle$, and $|9\rangle$ are even-parity, as evidenced by their finite overlap with $|0\rangle$ under the reflection-symmetric operator $\hat{J}_{x}^{2}$, see Fig. 7(b).
These symmetry considerations directly inform the structure of the photoexcitation pathways. Because $\hat{J}_{x}$ is odd under $\hat{\cal R}_{3}$, odd-parity excited states are directly accessible from the even-parity ground state via single-photon processes. In contrast, even-parity excited states cannot be reached at leading order and instead require higher-order processes involving even powers of $\hat{J}_{x}$, such as $\hat{J}_{x}^{2}$, which are reflection-symmetric. Consequently, at small pump amplitudes the dynamics is dominated by odd-parity states, while increasing the pump strength enables multiphoton processes that populate even-parity states through two-photon resonances and higher odd-parity states through three-photon pathways.
This structure explains the observed hierarchy of state populations under the drive. At small pump amplitudes, the dynamics is dominated by odd-parity states such as $|3\rangle$, which couple linearly to the even-parity ground state via single-photon processes mediated by $\hat{J}_{x}$. As the pump amplitude increases, higher-order processes become effective, allowing access to even-parity states such as $|2\rangle$ through two-photon resonances, as well as to higher odd-parity states such as $|6\rangle$ via three-photon pathways. The enhanced occupation of these states under a drive frequency $\Omega\simeq 1.713,t_{h}$ is therefore consistent with both their symmetry under reflection and their energetic proximity to the corresponding multiphoton resonance conditions.
We emphasize, however, that reflection symmetry alone does not determine the efficiency of state targeting. While symmetry dictates which matrix elements are allowed, the actual population transfer depends quantitatively on energy denominators, matrix-element magnitudes, and the pump amplitude. In particular, once the pump is applied, the time-dependent Hamiltonian explicitly breaks $\hat{\cal R}_{3}$, and parity ceases to be a conserved quantity during the dynamics. The observed state selectivity, therefore, reflects a combination of symmetry-based selection rules and dynamical resonance effects, both of which are explicitly captured in our numerical time-evolution and perturbative analyses.